Alternatives

Introduction

All of this is in the sipb locker unless otherwise stated. Type add sipb before using any of these.

Email

exmh An X Windows mail client more actively maintained than xmh. mh-rmail A way to read your mail in emacs. gnus Another featurefull way of reading your mail in emacs.

Zephyr

zion Replaces .zwgc.desc with perl and introduces feature based zephyr. vt An alternative to zwgc -ttymode. Xscretched This has no documentation.